# RLGNet: Repeating-Local-Global History Network for Temporal Knowledge Graph Reasoning

### v1

@@ 4.4 Global History Encoder
The global history encoder is designed to consider and extract information from a global perspective in order to capture global facts. For each query, we count the frequencies and the recent appearances of candidate entities. The module then uses an attention mechanism to integrate this information about the candidate entities. First, the attention scores for the time-related candidate entities of the current query are calculated:
$\mathcal{A}_{1}(q,i)=\sigma(W_{3}[h_{q};r_{q}])^{T}\sigma(W_{4}[\textbf{V}_{2}(t_{q}-t_{q}^{i});h_{i}])$ (13)
The attention scores for frequency-related candidates is similarly:
$\mathcal{A}_{2}(q,i)=\sigma(W_{5}[h_{q};r_{q}])^{T}\sigma(W_{6}[\textbf{V}_{3}(cnt^{i}_{q});h_{i}])$ (14)
@@
Considering the similarities between the characteristics of frequency and time, we employ a time vector to represent frequency. $W_{4}\in\mathbb{R}^{d\times 2d},W_{3}\in\mathbb{R}^{d\times 2d},W_{5}\in\mathbb{R}^{d\times 2d}$, and $W_{6}\in\mathbb{R}^{d\times 2d}$ are learnable parameters. $t_{q}^{i}$ and $cnt_{q}^{i}$ represent the frequency and last occurrence of candidate entity $i$ for query q. Additionally, taking computational resources into consideration, we only selected $top_{k}^{all}$ candidate entities. Following, we normalize their respective scores and multiply by the candidate entity representations to obtain representations $C_{q}^{gap}$ and $C_{q}^{cnt}$ for time-relevant and frequency-relevant candidates, respectively.
$C_{q}^{gap}=\sum_{i\in\mathcal{C}_{q}^{1:t_{q}-1}}\frac{\exp(\mathcal{A}_{1}(q,i))\cdot h_{i}}{\sum\limits_{j\in\mathcal{C}_{q}^{1:t_{q}-1}}\exp(\mathcal{A}_{1}(q,j))}$ (15)
$C_{q}^{cnt}=\sum_{i\in\mathcal{C}_{q}^{1:t_{q}-1}}\frac{\exp(\mathcal{A}_{2}(q,i))\cdot h_{i}}{\sum\limits_{j\in\mathcal{C}_{q}^{1:t_{q}-1}}\exp(\mathcal{A}_{2}(q,j))}$ (16)
The module then calculates the entity score for the global history encoder:
$score_{glo}^{q}=\varphi_{glo}(H,h_{q},r_{q},\textbf{V}_{4}(t_{q});C_{q}^{gap};C_{q}^{cnt})$ (17)
$h_{q}$ and $r_{q}$ represent the query entity and relation, Eq (17) can be viewed as an instance of Eq (3).
4.5 Repeating History Encoder
@@
The Repeating History Encoder is designed to improve prediction by increasing the weight of historical events. We first filter out and retain all the repetitive facts. Then, only the candidate entities of facts that rank in the top($top_{k}$) in terms of frequency are retained. In this module, the frequency of the candidate entities is also represented using a time vector. The repeating history encoder’s score function is a multilayer perceptron (MLP):
$\textbf{MLP}(x)=W_{7}\cdot\sigma(W_{8}\cdot\sigma(W_{9}\cdot x))$ (18)
$W_{7}\in\mathbb{R}^{1\times d}$, $W_{8}\in\mathbb{R}^{d\times 3d}$, and $W_{9}\in\mathbb{R}^{3d\times 5d}$ are learnable parameters. The extra score for candidates is calculated as follows:
$score_{rep}^{q,i}=\textbf{MLP}([h_{q};r_{q};\textbf{V}_{5}(t_{q});h_{i};\textbf{V}_{6}(cnt_{q}^{i})])$ (19)
It should be noted that when entity $i\notin\mathcal{C}_{q,top_{k}}^{1:t_{q}-1}$ the $score_{rep}^{q,i}$ is 0.
@@ 4.6 Loss Function
The model uses the cross-entropy loss function as a standard to measure loss, with the following formula:
$\mathcal{L}(score_{i},\mathcal{G}_{\mathcal{F}(i)})=\sum_{(s,r,o,t)\in\mathcal{G}_{\mathcal{F}(i)}}y_{t}\log P_{i}(o|s,r,t)$ (20)
where $P_{i}(o|s,r,t)=softmax(score_{i})$ represents the predicted probability of entities, and $y_{t}\in|\mathcal{E}|$is the label vector, where an element is 1 if the fact occurs, or 0 otherwise. Since each encoder in the model is trained separately, each one has a different loss function. When $i=rep$ , $\mathcal{F}(i)$ equals $rep$. And when $i\in\{loc,glo\}$, $\mathcal{F}(i)$ retrieves $raw$.
@@ 4.7 Final Score
The final model score combines scores from the repeating, local and global history encoders.
@@
Table 1: Performance (in percentage) on on ICEWS18, ICESW14, and ICEWS05-15.
@@ 5.1 Setup
ICE18 23,033 256 373,018 45,995 49,545 24 hours
ICE14 7,128 230 63,685 13,823 13,222 24 hours
ICE15 10,488 251 368,868 46,302 46,159 24 hours
@@
Datasets. We use six TKG datasets to evaluate the model’s effectiveness in entity prediction task, including ICEWS14(?), ICEWS18(?), and ICEWS05-15(?) from the Integrated Crisis Early Warning System (ICEWS) and the event-driven GDELT(?) dataset. Public datasets WIKI(?) and YAGO(?) are also included. All datasets are time-partitioned into Training (80%), Validation (10%), and Test (10%). More details on datasets are in Table 3.
Evaluation Metrics. To assess TKG reasoning performance, we used Mean Reciprocal Rank (MRR) and Hits@k metrics. MRR calculates average inverse rankings for actual entities across all queries, while Hits@k denotes the proportion of real entities appearing within the top k rankings. Previous studies have indicated that traditional filtering settings are flawed(?). Thus, we report the experimental results after using the time-aware filtering settings.
Baselines. RLGNet is compared with six baseline models on single-step reasoning task, including xERTE(?), RE-GCN(?), TITer(?), TiRGN(?), CEN(?), and RETIA(?). Since some models are not designed for multi-step reasoning task, we have selected a few models to report their performance on multi-step reasoning task, including CyGNet(?), RE-GCN(?), TiRGN(?), and CENET(?).
Implementation Details. For all datasets, we set the vector dimension $d$ to 200, and the time vector dimension $d_{2}$ in the local history encoder to 48. The $top_{k}$ and $top_{k}^{all}$ are set to 20 and 200. The number of GCN layers $\omega$ is set to 1. For ICEWS18, ICEWS14, ICEWS05-15, GDELT, WIKI, and YAGO, the adjacent history length $m$ is set to 10, 10, 15, 10, 1, and 1, respectively. The hyperparameter $\alpha$ is set to 0.8 in ICEWS, 0.9 in YAGO and WIKI, and 0.1 in GDELT. At the same time, a static graph constraint similar to that of RE-GCN is also added to ICEWS. Adam is used for parameter learning, with a learning rate set to 0.001. In the local history encoder, we employ StepLR to adjust the learning rate. In this setup, the value of gamma is set to 0.8. For the YAGO and WIKI datasets, the step-size value is set to 2, while for other datasets, step-size is 10. Please note that each encoder is trained independently, which results in them not sharing weights.
@@ 5.2 Results
The reasoning results of the entity prediction task are shown in Tables 1 and 2, where RLGNet outperforms other baselines in most cases. On six benchmark datasets, especially on the GDELT and ICEWS14 datasets, the MRR scores of RLGNet have increased by 3.42% and 2.11%, respectively. The primary reason for this improvement is that RLGNet not only considers repeating, local, and global historical information, but also integrates query information, which is not simultaneously considered by other models.
In single-step reasoning performance comparison, RLGNet demonstrates stronger performance in most cases. In the TKG extrapolation task, methods that consider query information usually achieve higher H@1 results, in comparison to the methods that ignore it(?). However, these methods may perform slightly worse on other metrics like H@3 and H@10. This could be because methods that don’t consider query information may introduce answers that don’t fully match the specific query, leading to inferior performance on H@1. This is also why, in the single-step reasoning task on the ICEWS05-15 dataset, even though TiRGN surpasses RLGNet by 0.08% in H@3 and 0.53% in H@10, RLGNet still outperforms TiRGN by 1.09% in H@1.
In multi-step reasoning task, the model is unable to know the ground truth facts during the test period. Local historical information relies on recently occurring facts, making it potentially noisier compared to single-step reasoning task. Local historical information relies on recently occurring facts, making it potentially noisier compared to single-step reasoning task. This is why global historical information becomes more important in multi-step reasoning task compared to single-step reasoning. TiRGN and RE-GCN has not fully considered global historical information, while CyGNet and CENET have overlooked local historical information. In contrast, RLGNet effectively integrates repetitive, local, and global historical information based on query information. Therefore, it achieves the best performance in multi-step reasoning task.
5.3 Ablation Study
Model ICEWS18 ICWS14 ICEWS05-15 WIKI YAGO GDELT
@@ 5.3 Ablation Study
Table 4: The MRR (in percentage) results of the ablation studies.
$top_{k}$ ICE18 ICE14 ICE15 WIKI YAGO GDELT 5 34.49 51.94 42.77 26.69 84.05 77.67 10 40.29 58.26 47.10 34.26 90.76 85.37 20 44.63 62.90 49.82 42.35 92.63 86.96 30 46.58 64.80 50.69 47.05 92.73 87.03 100 49.76 67.59 51.75 58.41 92.73 87.04 $\infty$ 50.42 68.38 51.85 64.93 92.73 87.04
Table 5: The proportion of repeating facts when $top_{k}$ is set to different values.
To validate the impact of different modules, we conducted ablation experiments on single-step and multi-step reasoning tasks on six benchmark datasets. The results are shown in Tables 4. Our ablation studies indicated that the local history encoder (Loc) and the global history encoder (Glo) do not always improve results. Specifically, the use of Glo and Loc modules led to a decrease in performance in single-step reasoning tasks in the ICEWS05-15 and GDELT datasets. However, considering that their improvement in other tasks far outweighs the decline in these two subtasks, we thus believe Glo and Loc are beneficial. On the other hand, the Repetitive History Encoder (Rep) can improve the final performance across various tasks, suggesting that events usually exhibit a degree of repetition. As shown in Table 5, the proportion of repetitive events in the dataset well corroborates this point. Overall, while Loc and Glo may not perform ideally in some specific tasks, these three historical information encoding methods can all contribute to better prediction results.
@@ 5.4 Hyperparameter Analysis
Figure 3: The impact of $\alpha$ on MRR (in percentage) results.
Effect of $\alpha$ Values in RLGNet. When processing multi-step and single-step reasoning tasks, global and local encoders contribute differently to the answers, as shown in Figure 3. In single-step reasoning task, we find that on the ICEWS dataset, the performance is best when $\alpha$ is between 0.6 and 0.8. On the WIKI and YAGO datasets, the performance peaks when $\alpha$ is between 0.8 and 1. On the GDELT dataset, the optimum range for $\alpha$ is between 0 and 0.2. Therefore, we set $\alpha$ to 0.8 and 0.1 in ICEWS and GDELT respectively, and to 0.9 in both WIKI and YAGO. In multi-step reasoning task, in order to achieve optimal performance, the value of $\alpha$ is usually smaller than the value of $\alpha$ in single-step inference tasks. This is because the local history encoder can more easily capture the sequential pattern of facts when the baseline truth is known. It’s noteworthy that in the GDELT dataset, the contribution of global historical information always exceeds that of local historical information. This is due to the GDELT dataset containing more abstract entity concepts, making it more difficult for the model to capture the sequential patterns of facts(?).

### v2

@@ 4.3 Global History Module
The Global History Module is designed to consider and extract information from a global perspective in order to capture global facts. For each query, we calculate the occurrence frequency of candidate entities and the last time they appeared. The module then uses an attention mechanism to integrate this information about the candidate entities.
$\mathcal{A}_{1}(q,i)=(W_{1}^{G}[h_{q}^{G};r_{q}^{G}])^{T}(W_{2}^{G}[\textbf{V}_{1}^{G}(t_{q}-t_{q}^{i});h_{i}^{G}])$ (12)
The attention scores for frequency-related candidates are calculated similarly:
$\mathcal{A}_{2}(q,i)=(W_{3}^{G}[h_{q}^{G};r_{q}^{G}])^{T}(W_{4}^{G}[\textbf{V}_{2}^{G}(cnt^{i}_{q});h_{i}^{G}])$ (13)
@@
where $h_{q}^{G}$, $r_{q}^{G}$, and $h_{i}^{G}$ refer to the initial embeddings of the query entity, relation, and candidate entity $i$ in the Global History Module. $\textbf{V}_{1}^{G}$ and $\textbf{V}_{2}^{G}$ are Numerical Embeddings. $W_{1}^{G}$, $W_{2}^{G}$, $W_{3}^{G}$, and $W_{4}^{G}$ are learnable parameters. $t_{q}^{i}$ and $cnt_{q}^{i}$ represent the last occurrence and frequency of candidate entity $i$ for query $q$. Additionally, taking computational resources into consideration, we only selected the $top_{k}^{all}$ candidate entities. Following this, we normalize their respective scores and multiply by the candidate entity embeddings to obtain the embedding $C_{q}^{gap}$ for time-relevant:
$C_{q}^{gap}=\sum_{i\in\mathcal{C}_{q,top_{k}^{all}}^{1:t_{q}-1}}\frac{\exp(\mathcal{A}_{1}(q,i))\cdot h_{i}^{G}}{\sum\limits_{j\in\mathcal{C}_{q,top_{k}^{all}}^{1:t_{q}-1}}\exp(\mathcal{A}_{1}(q,j))}$ (14)
We normalize $\mathcal{A}_{2}$ and multiply it by the entity embeddings, then sum the results to obtain the frequency-related candidate embedding $C_{q}^{cnt}$, just as we do with $C_{q}^{gap}$. Similar to Equation 11, the score for the Global History Module is:
$score_{Glo}^{q}=H_{G}\cdot\phi_{G}(h_{q}^{G},r_{q}^{G},\textbf{V}_{3}^{G}(t_{q}),C_{q}^{gap},C_{q}^{cnt})$ (15)
where $H_{G}$ represents the initial embeddings of all entities in the Global History Module. $\textbf{V}_{3}^{G}$ is Numerical Embeddings. $\phi_{G}$ represents the score decoder based on ConvTransE in the Global History Module.
4.4 Repeating History Module
@@
The Repeating History Module is designed to improve prediction by increasing the weight of historical events. We first filter and retain the repeated facts, and refer to the collection of these repeated facts as $\mathcal{G}_{rep}$. Then, only the candidate entities of facts that rank in the top $top_{k}$ in terms of frequency are retained.
The Repeating History Module’s score function is a three-layer MLP. The candidate entities’ scores are calculated using the MLP as follows:
$score_{Rep}^{q,i}=MLP([h_{q}^{R};r_{q}^{R};\textbf{V}_{1}^{R}(t_{q});h_{i}^{R};\textbf{V}_{2}^{R}(cnt_{q}^{i})])$ (16)
where $h_{q}^{R}$,$h_{i}^{R}$ and $r_{q}^{R}$ represent the initial embeddings of the query entity, candidate entity, and relation in the Repeating History Module, respectively. $\textbf{V}_{1}^{R}$ and $\textbf{V}_{2}^{R}$ are Numerical Embeddings. It should be noted that when candidate entity $i\notin\mathcal{C}_{q,top_{k}}^{1:t_{q}-1}$, the $score_{Rep}^{q,i}$ is 0.
@@ 4.5 Loss Function
Since the model employs an ensemble learning strategy, each module is trained independently and uses the cross-entropy loss function to calculate the loss value:
$\mathcal{L}(score,\mathcal{G})=\sum_{\mathcal{G}^{t}\in\mathcal{G}}\sum_{(s,r,o,t)\in\mathcal{G}^{t}}y_{t}\log P(o|s,r,t)$ (17)
The score represents the predicted scores of each module, and $\mathcal{G}$ represents the sequence of snapshot graphs. Here, $P(o|s,r,t)=\text{softmax}(score)$ represents the predicted probability of entities, and $y_{t}\in|\mathcal{E}|$ is the label vector, where an element is 1 if the fact occurs, or 0 otherwise. Therefore, the loss functions for the three modules are $\mathcal{L}(score_{Loc},\mathcal{G}_{raw})$, $\mathcal{L}(score_{Glo},\mathcal{G}_{raw})$, and $\mathcal{L}(score_{Rep},\mathcal{G}_{rep})$, respectively.
@@ 4.6 Final Score
The final model score combines scores from the Repeating, Local, and Global History Modules:
@@
Table 1: Performance (in percentage) on ICEWS18, ICEWS14, and ICEWS05-15.
@@ 5 Experiments
ICEWS18 23,033 256 373,018 45,995 49,545 24 hours
ICEWS14 7,128 230 63,685 13,823 13,222 24 hours
ICEWS05-15 10,488 251 368,868 46,302 46,159 24 hours
@@
5.1 Setup
Datasets
We use six TKG datasets to evaluate the model’s effectiveness in the entity prediction task, including ICEWS14 (?), ICEWS18 (?), and ICEWS05-15 (?) from the Integrated Crisis Early Warning System (ICEWS), and the event-driven GDELT (?) dataset. Public datasets WIKI (?) and YAGO (?) are also included. All datasets are time-partitioned into Training (80%), Validation (10%), and Test (10%). More details on the datasets are provided in Table 3.
Evaluation Metrics
To assess TKG reasoning performance, we use Mean Reciprocal Rank (MRR) and Hits@k metrics. MRR calculates the average inverse rankings for actual entities across all queries, while Hits@k denotes the proportion of real entities appearing within the top k rankings. Therefore, higher values of MRR and Hits@k indicate better model performance. Previous studies have indicated that traditional filtering settings are flawed (?). Thus, we report the experimental results after using the time-aware filtering settings.
Baselines
RLGNet is compared with six baseline models on the single-step reasoning task, including xERTE (?), RE-GCN (?), TITer (?), TiRGN (?), CEN (?), and RETIA (?). Since some models are not designed for the multi-step reasoning task, we have selected a few models to report their performance on the multi-step reasoning task, including CyGNet (?), RE-GCN (?), TiRGN (?), and CENET (?).
Implementation Details
For all datasets, we set the embedding dimension to 200, and the dimension of $\textbf{V}_{1}^{L}$ in the Local History Module to 48. The $top_{k}$ and $top_{k}^{all}$ are set to 20 and 200, respectively. The number of GCN layers $\omega$ is set to 1. For ICEWS18, ICEWS14, ICEWS05-15, GDELT, WIKI, and YAGO, the adjacent history length $m$ is set to 10, 10, 15, 10, 1, and 1, respectively. The hyperparameter $\alpha$ is set to 0.8 in ICEWS, 0.9 in YAGO and WIKI, and 0.1 in GDELT. Additionally, a static graph constraint similar to that of RE-GCN is also added to ICEWS. Adam is used for parameter learning, with a learning rate set to 0.001. In the Local History Module, we employ StepLR to adjust the learning rate. In this setup, the learning rate decay is set to 0.8. For the YAGO and WIKI datasets, the step-size value is set to 2, while for other datasets, the step-size is 10. Please note that each module is trained independently, which results in them not sharing weights.
@@ 5.2 Results
The reasoning results of the entity prediction task are shown in Tables 1 and 2, where RLGNet outperforms other baselines in most cases.On six benchmark datasets, RLGNet has demonstrated superior performance in both multi-step and single-step reasoning tasks in most cases, particularly on the GDELT and ICEWS14 datasets, where its MRR scores have increased by 3.42% and 2.11%, respectively. This improvement is primarily due to RLGNet’s hybrid architecture and ensemble learning design. Each independent module considers historical information from different scales and constrains the answer range using candidate entities. The hybrid architecture design enhances the model’s performance in multi-task scenarios, ultimately leading to more accurate predictions. Compared to various single architecture models such as RE-GCN and TiRGCN, RLGNet exhibits stronger generalization capabilities in both multi-step and single-step reasoning. Moreover, compared to models that do not consider multiple scales, such as CEN and CyGNet, RLGNet shows superior performance.
However, it is worth noting that in the single-step reasoning task on the ICEWS05-15 dataset, RLGNet’s H@3 and H@10 metrics are 0.08% and 0.53% lower than those of TiRGCN, respectively, but its H@1 metric is 1.09% higher than that of TiRGCN. We believe this phenomenon is because, in TKG extrapolation tasks, methods that consider candidate entities typically achieve higher H@1 results compared to those that ignore candidate entities. However, these methods may perform slightly worse on other metrics such as H@3 and H@10. This is because introducing candidate entities constrains the answer range, but candidate entities are sampled based on past facts, making the model more inclined towards past patterns. As a result, the model may overlook new emerging patterns or changes when predicting the future. Therefore, although candidate entities can improve the accuracy of the H@1 metric, they may lack diversity and breadth.
Model ICEWS18 ICEWS14 ICEWS05-15 WIKI YAGO GDELT
@@
Table 4: The MRR (in percentage) results of the ablation studies. Single and Multi represent single-step and multi-step reasoning, respectively. Since Rep can only predict repeating facts, we did not report the MRR score for Rep.
$top_{k}$ ICEWS18 ICEWS14 ICEWS05-15 WIKI YAGO GDELT
5 34.49 51.94 42.77 26.69 84.05 77.67
10 40.29 58.26 47.10 34.26 90.76 85.37
20 44.63 62.90 49.82 42.35 92.63 86.96
30 46.58 64.80 50.69 47.05 92.73 87.03
100 49.76 67.59 51.75 58.41 92.73 87.04
$\infty$ 50.42 68.38 51.85 64.93 92.73 87.04
Table 5: The proportion (in percentage) of repeating facts when $top_{k}$ is set to different values.
5.3 Ablation Study
To validate the impact of different modules, we conducted ablation experiments on single-step and multi-step reasoning tasks across six benchmark datasets. The results are shown in Table 4. Our ablation experiments studied the impact of the Repeating History Module (Rep), Local History Module (Loc), and Global History Module (Glo) on performance. From the hybrid architecture perspective, the Loc based on the RNN architecture and the Glo based on the MLP architecture can complement each other in capability. For instance, Glo+Loc always outperforms the individual Loc or Glo modules. From the multi-scale perspective, information at different scales is generally beneficial to the model’s performance. Analyzing from the perspective of repeating history information scale, we found that adding Rep to Loc, Glo, or Loc+Glo always improves the model’s performance. We believe the enhancement brought by repeating history information is due to the significant proportion of repeating facts in TKG datasets, which is evident from Table 5, where the proportion of repeating facts in the YAGO dataset reaches as high as 92.73% when $top_{k}$ is set to $\infty$.
Regarding global and local history information scales, their performance varies due to differences between datasets (specific details can be found in Table 2). Since local history information only considers facts within a unit time period, the richness of facts occurring within this period determines the performance of Loc. For datasets with large time spans, such as the WIKI and YAGO datasets where the unit time is one year, all facts occurring within the same year are included in one time unit. Therefore, each unit time contains more and richer facts, which is why Loc performs better. However, for datasets with smaller time spans, such as the GDELT dataset where the unit time is 15 minutes, Glo performs better because it can better grasp the characteristics of the entire dataset.
Overall, although in single-step reasoning tasks on the ICEWS05-15 and GDELT datasets, using Glo+Loc+Rep resulted in lower performance than Loc+Rep and Glo+Rep, the improvements in other tasks far exceeded the declines in these two sub-tasks. Therefore, we believe that the hybrid multi-scale architecture is effective in TKG reasoning tasks.
Figure 3: The impact of $\alpha$ on MRR (in percentage) results.
@@ 5.4 Hyperparameter Analysis
We investigated the impact of the values of $\alpha$, $\omega$, and $top_{k}$ on the performance of RLGNet. The contributions of the Global and Local History Modules to the answers vary in single-step and multi-step reasoning tasks, which can be observed from the value of $\alpha$, as shown in Figure 3. In single-step reasoning tasks, we found that on the ICEWS dataset, the performance is optimal when $\alpha$ is between 0.6 and 0.8. On the WIKI and YAGO datasets, the performance peaks when $\alpha$ is between 0.8 and 1. On the GDELT dataset, the optimal range for $\alpha$ is between 0 and 0.2. Therefore, we set $\alpha$ to 0.8 and 0.1 for ICEWS and GDELT, respectively, and 0.9 for WIKI and YAGO.
In multi-step reasoning tasks, the value of $\alpha$ usually needs to be smaller than in single-step reasoning tasks to achieve optimal performance. This indicates that the Global History Module based on MLP is more important in multi-step reasoning, whereas the Local History Module based on RNN is more crucial in single-step reasoning. The experimental results also demonstrate the unique advantages of the methods based on MLP and RNN architectures, validating our hybrid architecture design’s ability to leverage the strengths of both.
Additionally, we studied the impact of $\omega$ and $top_{k}$ values on the performance of RLGNet on the ICEWS14 and YAGO datasets, as shown in Figure 4. As $\omega$ increases, the performance of the model on ICEWS14 fluctuates slightly, while the performance on the YAGO dataset rapidly declines. Therefore, we set $\omega$ to 1 for all datasets. In the YAGO dataset, setting $top_{k}$ at a higher value yields better performance; however, changes in $top_{k}$ do not significantly affect the performance on the ICEWS14 dataset. Therefore, we set $top_{k}$ for all datasets to 20.
5.5 Effectiveness of Ensemble Learning Strategy
To demonstrate the effectiveness of the ensemble learning strategy, we compared the performance of different modules using ensemble learning and joint learning strategies on the YAGO and ICEWS14 datasets, as shown in Table 6. In joint learning, all modules share the embeddings of entities and relationships, with other settings being the same as those in the ensemble learning strategy. We found that the performance of any module combination using the ensemble learning strategy was generally superior to that of the joint learning strategy. This is because, in TKG, the patterns of facts change over time, making it inevitable that there will be noisy data in the historical information. Although joint learning, through weight sharing, allows the model to perform better during the training phase, it also leads to fitting the noisy data, thereby impairing the model’s ability to predict future events. Therefore, we believe that the ensemble learning strategy is an effective approach for inference tasks in TKGs.